OpenStudy (anonymous):

When a certain spring is expanded 6 centimeters from its natural position and held fixed, the force necessary to hold it is 1200 dynes. Find the work W required to stretch the spring an additional 6 centimeters.

OpenStudy (anonymous):

What's the spring constant?

OpenStudy (anonymous):

Never mind you can calculate it with the given information

OpenStudy (anonymous):

I tried \[\int\limits_{0}^{6} (k x)dx=1200 \] \[k*(6^2/2)=1200 \] So that k=200/3

OpenStudy (anonymous):

The force on a spring is F= k (d1 - d0) we say d0=0 and d1 is 6 cm so 1200= k * 6 200= k F= k ( d1- d0) where d1 is 12 cm and d0 is 6 cm F to stretch additional 6 cm is 200* 6 F= 1200 where work is F* change in distance we changed the distance by 6 cm so it should be 1200 * 6= 7200

OpenStudy (anonymous):

I solved this with the knowledge I have in physics so this might not be what your teacher wants to see.

OpenStudy (anonymous):

Thank you I'll try it out

OpenStudy (anonymous):

Physics is a useful tool in mathematics: it explains a lot; math concentrates on the calculations.